6x^2+5y=58
What are/is the intercept(s)?

Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

If x = 0
5y = 58
y = 58/5
y-intercept is: (0,58/5)
If y = 0
6x^2=58
x^2 = 58/6 = 29/3
x = sqrt(29/3) or x = -sqrt(29/3)
x-intercepts are: (-sqrt(29/3),0) and (sqrt(29/3),0)

If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks

Answer by stanbon(75887)

6x^2+5y=58
------
y-intercept ??
Let x = 0, then y = 58/5 = 11 3/5
----------------
x-intercepts ??
Let y = 0, then x = +-sqrt(58/6) = +- sqrt(29/3)
================
Cheers,
Stan H.
==================